OpenStudy (anonymous):

Can someone help me.. If f(x)=3x^2-x^3, find f'(1)

OpenStudy (anonymous):

6x-3x^2

OpenStudy (anonymous):

How'd you get that...

OpenStudy (anonymous):

now plug in the one to get 6(1)-3(1)^2

OpenStudy (anonymous):

oh are you just starting calculus I?

OpenStudy (anonymous):

lol, yes...

OpenStudy (anonymous):

then you probably havnt learned the shortcut

OpenStudy (anonymous):

you used the power rule right?

OpenStudy (anonymous):

nx^n-1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

just bring the exponent down and multiply it by whats in front of x, then subtract 1 from numerater

OpenStudy (anonymous):

But, my professor wants us to do it the longer way:/ or we won't get credit.. can you help me using that way? f(a+h)-f(a)/h

OpenStudy (anonymous):

yes i will one sec

OpenStudy (anonymous):

so lets say you have f(x)=x^2

OpenStudy (anonymous):

and you want to find f'(3)

OpenStudy (anonymous):

so u start with \[\lim_{h \rightarrow infinity }\] of that equation you posted

OpenStudy (anonymous):

\[f(x)=3x^2-x^3\]To find the derivative, we know that the derivative of a sum or difference is equivalent to the sum or difference of the derivatives of the terms.\[f'(x) = \frac{d(3x^2)}{dx}-\frac{d(x^3)}{dx}\\\ \ \ \ \ \ \ \ =6x-3x^2\]Now, evaluate the derivative for 1.\[f'(1)=6(1)-3(1)^2=6-3=3\]

OpenStudy (anonymous):

does that help

OpenStudy (anonymous):

if you have not got to the power rule yet, and you just want \(f'(1)\) you can use \[\lim_{x\to 1}\frac{f(x)-f(1)}{x-1}\] or \[\lim_{h\to 0}\frac{f(1+h)-f(1)}{h}\] not sure what method you were taught

OpenStudy (anonymous):

sometimes at the beginning it is easier to work directly with numbers

OpenStudy (anonymous):

thanks guys :) you're so much help!

OpenStudy (anonymous):

did i get a medal or whatever lol

OpenStudy (anonymous):

\[f(x)=3x^2-x^3\] \[f(1)=3-1=2\] \[\lim_{x\to 1}\frac{f(x)-f(1)}{x-1}\] \[=\lim_{x\to 1}\frac{3x^2-x^3-2}{x-1}\] \[=\lim_{x\to 1}\frac{-(x-1) (x^2-2 x-2)}{x-1}\] \[=\lim_{x\to 1}-(x^2-2x-2)\]

OpenStudy (anonymous):

power rule is undoubtedly easier but if you have to do it by hand you can write this

OpenStudy (anonymous):

For the more rigorous proof...\[\lim_{\delta{x}\to\infty} \frac{f(x+\delta{x})-f(x)}{\delta{x}}\\=\lim_{\delta{x}\to\infty} \frac{3(x+\delta{x})^2-(x+\delta{x})^3-3x^2+x^3}{\delta{x}}\\=\lim_{\delta{x}\to\infty} \frac{3(x^2+2x\delta{x} + {\delta{x}}^2)-x^3-3x^2\delta{x}-3x{\delta{x}}^2-3x^2+x^3}{\delta{x}}\\=\lim_{\delta{x}\to\infty} \frac{3x^2+6x\delta{x}+3{\delta{x}}^2 - x^3 - 3x^2\delta{x} - 3x\delta{x}^2-3x^2+x^3}{\delta{x}}\\=\lim_{\delta{x}\to\infty} \frac{6x\delta{x} +3\delta{x}^2-3x^2\delta{x}-3x\delta{x}^2}{\delta{x}}\\=\lim_{\delta{x}\to\infty}[6x+3\delta{x}-3x^2-3x\delta{x}^]\\=6x-3x^2\]